OpenStudy (lgbasallote):

Find the general solution: \[(3+y+2y^2\sin^2 x)dx + (x+ 2xy -y\sin 2x)dy=0\]

OpenStudy (lgbasallote):

\[\frac{\partial M}{\partial y} = 1 + 4y\sin^2 x\] \[\frac{\partial N}{\partial x} = 1 + 2y - 2y\cos 2x\] so i guess tis isnt exact? or did i do something wrong?

OpenStudy (anonymous):

wait wat's 1-cos2x?

OpenStudy (anonymous):

it's 2sin^2x......it's exact

OpenStudy (lgbasallote):

wait what?

OpenStudy (anonymous):

u kno 1-cos2x = 2sinx*sinx

OpenStudy (lgbasallote):

yeah?

OpenStudy (lgbasallote):

oh i see your point

OpenStudy (anonymous):

actually um wateva u call it it's root((1-cos2x)/2) = sinx

OpenStudy (lgbasallote):

so it's exact..

OpenStudy (lgbasallote):

\[\int (3+y+2y^2 \sin^2 x)dx\] \[3x + xy + y^2(x - \sin x \cos x)\]

OpenStudy (lgbasallote):

then i take derivative of y

OpenStudy (lgbasallote):

\[x + 2yx - 2y\sin x\cos x + g'(y) = x+2xy - y\sin 2x\] \[g'(y) = 2y\sin x \cos x - y\sin 2x\] \[g(y) = 2\sin x \cos x - \sin 2x\] so the G.S. would be \[3x + xy + y^2(x - \sin x \cos x) + 2\sin x \cos x - \sin 2x + C\] correct?

OpenStudy (anonymous):

u kno there's a direct solution for exact equations

OpenStudy (lgbasallote):

really? what?

OpenStudy (anonymous):

i dont remember xactly but it's smthin llike integrate m(x,y) keeping y/x as a constant = integrate n(x,y) not containing terms of x/y + c

OpenStudy (lgbasallote):

oh that..yeah...im more comfortable with this though

OpenStudy (lgbasallote):

im more comfortable when the solution is one thread instead of inconsistent

OpenStudy (anonymous):

ok